OpenStudy (anonymous):

graph each equation. identify the x and y-intercepts and the asymptotes of the graph. y=-1(over) x -4

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Okay, you should be able to do this one. It's like the one I drew for you as "extra credit", except shifted by 2 more units along the x axis, and the flipped over (because it is -1/ instead of 1/)

OpenStudy (anonymous):

it never hits the line again. so is when the denominator cant be 0 and x=2

OpenStudy (whpalmer4):

Let me check something: the function is \[y = \frac{-1}{x-4}\]Right? Not \[y = \frac{-1}{x} - 4\]

OpenStudy (anonymous):

no it is the second one

OpenStudy (whpalmer4):

Okay, glad I asked! That's a different transformation entirely!

OpenStudy (anonymous):

ok good i was like that doesn't look like the one you gave me :)

OpenStudy (whpalmer4):

Let's build this up piece by piece. Here's the graph of \(1/x\):

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

continue

OpenStudy (whpalmer4):

Now we have \(-1/x\), so wherever this one is positive, the other one is negative, and vice versa. In other words, we flip it over the \(x\) axis.

OpenStudy (whpalmer4):

Looks like this:

OpenStudy (anonymous):

ok i have the equation for the problem in the question graphed on my calculator

OpenStudy (whpalmer4):

Now that was one part of the transformation. The next part is that we subtracted 4 from it. If we subtract 4 from the result of the function, that just moves the whole graph down by 4 units, right?

OpenStudy (anonymous):

yes!!!

OpenStudy (whpalmer4):

So that also moves the horizontal asymptote down by 4. Are we agreed that the asymptotes of \(1/x\) and \(-1/x\) are identical?

OpenStudy (whpalmer4):

Please say "yes, of course!" :-)

OpenStudy (anonymous):

i think they are. they are identical right?

OpenStudy (anonymous):

ok i was right

OpenStudy (whpalmer4):

Right. So one asymptote is going to stay the same, and one is going to change when we shift the whole graph down by 4 units. Can you tell me which is which?

OpenStudy (anonymous):

so they asymptotes of -1/x is also x=0 and y=0 right

OpenStudy (anonymous):

the graph will change the x=0 right? they y=0 will stay the same

OpenStudy (whpalmer4):

Yes, \(-1/x\) has asymptotes \(x = 0\) (vertical) and \(y = 0\) (horizontal)

OpenStudy (anonymous):

ok is my other answer correct

OpenStudy (whpalmer4):

Okay, I've updated my graph to include the line y = -4

OpenStudy (anonymous):

no x=0 stays the same and the y will change!! Right????

OpenStudy (whpalmer4):

yes, that's correct!

OpenStudy (anonymous):

yes!!!!!! ok so will it change to y=4 or what??

OpenStudy (anonymous):

also will the x and y intercepts still be 0

OpenStudy (whpalmer4):

uh, here's my graph with the line y = 4 added...

OpenStudy (anonymous):

oh its -4 sorry :/

OpenStudy (whpalmer4):

remember, we had asymptotes of \(x = 0\) and \(y = 0\) before we shifted everything down by 4. when we shifted everything down by 4, the asymptotes also shifted, so the horizontal asymptote became \(y = -4\). The vertical asymptote is an infinitely long line, so we can't tell any different when it shifts :-)

OpenStudy (whpalmer4):

Again, let's solve for the intercepts: put in \(x=0\) and solve for \(y\). Then put in \(y = 0\) and solve for \(x\).

OpenStudy (whpalmer4):

You'll get one answer you've seen before, and one you haven't.

OpenStudy (whpalmer4):

(that should have been "The vertical asymptote is an infinitely long line, so we can't tell any different when it shifts VERTICALLY")

OpenStudy (anonymous):

ok so the asymptotes ar y=-4 and x=0 and the intercepts are y=0 and x=-5??

OpenStudy (whpalmer4):

Wait a minute, how did you get those values for the intercepts? \[y = -\frac{1}{x}-4\]\[y = -\frac{1}{0} -4 = \]

OpenStudy (whpalmer4):

And \[0 = -\frac{1}x - 4\]\[4 = -\frac{1}{x}\]\[4x=-1\]\[x =\]

OpenStudy (anonymous):

well when i plugged in 0 for x it said there was a error so i thought that y=0 oh and the x will be x=-1/4 or -.25

OpenStudy (whpalmer4):

Yeah, if there's an error, there's no value. So the correct answers are "no y-intercept" and "x = -1/4" and if you look at my graphs, you'll see the x-axis gets crossed at \(x=-1/4\) and the y-axis isn't crossed (because there's a vertical asymptote there).

OpenStudy (anonymous):

ok so the intercepts are no y-intercept and x=-.25. the asymptotes are x=0 and y=-4 correct?

OpenStudy (whpalmer4):

yes! spoken like a true math geek :-)

OpenStudy (whpalmer4):

though I'd go with x = -1/4 personally

OpenStudy (anonymous):

lol ok thank you so much. i understand asymptotes now also. you are a lifesaver!!!! thank you again

OpenStudy (whpalmer4):

it comes out exactly here, but not all fractions do. 1/3 = 0.333333333333...

OpenStudy (whpalmer4):

not to rain on your parade, but you can also have asymptotes that aren't parallel to either x or y axis, and they are trickier to figure out :-) probably don't need to know about them yet, though!

OpenStudy (whpalmer4):

at least I hope not, because I don't feel up to doing a good explanation of them :-)

OpenStudy (anonymous):

i dont think so :D

OpenStudy (whpalmer4):

well, if you need a refresher in a few days on this stuff, you know where to find me. I'm not always online, but I'll respond when I am.

OpenStudy (anonymous):

ok thank you